the youth organization has collected nickles and dimes for a charity drive.by the end of day 1, the children had $32.50. if there was two times as many dimes as nickels, how many of each coin was collected?

Answer by checkley79(3341)

LET X= THE NUMBER OF NICKLES.
LET 2X = THE NUMBER OF DIMES.
.05X+.10*2X=32.50
.05X+.20X=32.50
.25X=32.50
X=32.50/.25
X=130 NUMBER OF NICKLES.
2*130=260 NUMBER OF DIMES.
.05*130+.10*260=32,50
6.50+26.00=32.50
32.50=32.50